[Halld-jlab] Hall D Staff Meeting Notes, July 1, 2013

Mark M. Ito marki at jlab.org
Wed Jul 3 10:24:05 EDT 2013


Folks,

Find the minutes at 
https://halldweb1.jlab.org/wiki/index.php/Private:Hall_D_Staff_Meeting,_July_1,_2013 
and below.

   -- Mark
________________________________________

Private:Hall D Staff Meeting, July 1, 2013

    Present: Fernando Barbosa, Vladimir Berdnikov, Eugene Chudakov (chair),
    Hovanes Egiyan, Nerses Gevorgyan, Mark Ito, Lubomir Pentchev, Sergey
    Pozdniakov, Elton Smith, Steve Smith, Sascha Somov, Simon Taylor,
    Elliott Wolin, Beni Zihlmann

Steve

      * showed slides on analysis of reported events over the past year
        and statistics from observations by safety teams in visits made
        around the lab site

https://halldweb1.jlab.org/talks/2013-3Q/Road%20Show%20trend%20analysis%20June%202013.pptx

      * showed slides summarizing results of the periodic safety
        culture surveys
           + areas for improvement highlighted
           + specific written comments were presented

https://halldweb1.jlab.org/talks/2013-3Q/Road%20Show%20DSC%20Safety%20Culture%206-13-2013%20r1.pptx

Eugene

      * on schedule now, but lots of activities coming up with associated
        schedule risk
           + will have to progress at 1.2 M$ per month for some months
           + already spent 300 k$ on first three months of rebaselined
             schedule, about 10%
      * solenoid status
           + will have to start the refrigerator again soon
           + this will be last run before major surgery to refrigerator
                o carbon dust due to deteriorated charcoal filter, dust
                  clogs the plumbing
                o surgery involves bypassing the filter
                o similar procedure tried at Argonne with success
                o several months to complete
           + lifetime with carbon problem about 2 weeks
           + power supply still not ready
                o see transients in voltages which prevents low-threshold
                  in the quench detector
                o bug in firmware also discovered, new eprom required to
                  fix, probably will not be done
                o power supply tuned to 2 H, not 25 H, retuning being
                  considered, requires adjustment to electronic
                  configuration
           + a lot of discussion going on about current limit for the
             high-current test
           + refrigerator cool-down schedule for Sat. July 6

Beni

      * working on FDC read-out test on a full package
      * some noise problems when last set of channels connected
      * problem with FADCs different versions of firmware, different
        effective thresholds were being set
      * reliability problem seen with connections on one pre-amp
      * noise looks good now
      * analyzing efficiency of cathode strips

Mark

      * new complete build scripts deployed, feedback from a couple of
        users so far
      * start counter counters from McNeal received
           + five of them
           + look good visually
           + tested: good timing resolution, best seen so far
      * FADCs sent to FSU for TOF testing
           + Sasha has been able to read-out the flash
           + one board had a blown fuse, second board was sent, fuse blown
             again
           + modules have been received back at JLab, will be repaired and
             one sent back to FSU
           + may have to send a power supply

Elton

      * BCAL construction
           + 7 modules to go
           + enough light guides for all but 3
           + shipped non-conforming light guides to Santa Maria, new ones
             also shipped back
           + students: Laura and Tegan going back to IU this week
           + Tina will be back next week, will probably not train new
             students on construction
           + John Leckey will get new students to help on cabling the FCAL
      * EE from Santa Maria, Alan Toro, will be here for six months
        starting in July
           + will work on monitoring fan-out boards
           + experience at CERN, noise reduction
      * pion polarizability proposal approved
           + thinking of a workshop on August 23
                o Barbara Pasquini may visit
                o invite new collaborators
                o planning for a real design

Elliott

      * having a monitoring data challenge in late summer
           + David L. working on online jana application, translation
             tables
           + Sean Dobbs working on RootSpy
      * David Abbott is back and will start working on on high rate testing
      * ROC bids cam in under budget, Concurrent will be the vendor
      * using C++11 standard, getting experience now
      * monitoring plug-ins written, some interest from detector experts
      * inventory system progressing
      * software for quench protection being developed
      * hardware quench protector is being tuned, dials added
      * cabling planning in progress, Tim wants comprehensive list

Simon

      * making progress on FDC geometry
      * also working on bug fixes and improvements for the next data
        challenge

Sascha

      * working with Liping's students on fabrication of counters
           + checking PMTs
           + using UV glue for light guide to PMT joint
           + checking with cosmics
      * progress on trigger test set-up
           + synchronized 3 crates
           + loaded playback mode
           + added GTP
           + trying to understand the registers
           + do not have libraries for many boards
           + work on trigger module scalers
           + work on hit pattern trigger with microscope
      * fine-granularity pair spectrometer counters
           + all 1 mm counters done
           + 70 of 200 2 mm counters done

Lubomir

      * looking at problem with FDC pre-amp, flaky connection
      * next test: add F1 TDCs to read-out
      * at Blue Crab spare package
           + needed gusset rings temporarily for spacer
           + in process of re-install on spare package, non-flatness
             introduced, deformed by 1 cm
           + able to fix
      * thick cell on spare
           + able to see single electrons drifting in to wires
           + able to measure distribution of avalanche size for single
             electrons
           + using 40/60 Ar/CO[2]

Hovanes

      * desktops installed in counting house
      * cubicles ready to use
      * SOW for server procurement written: includes HPC nodes, our nodes,
        and infiniband switches, PR not submitted yet
      * Facilities issued PR for fiber between tagger service building and
        tagger hall
           + in August can start installing controls
      * working on progress reports and re-baseline tweaking
           + noticed that RF signal hardware not in project
           + need 2 boxes, one in Hall, one in tagger hall
      * work on EPICS alarms
           + using CSS
           + problem with user authentication
           + computer center setting up LDAP server for us
      * goniometer test
           + Vanik and Yi working on this
           + needed new build of EPICS base

Nerses

      * CAEN HV control drivers close to finished
           + fixing bugs
           + next: gui

Sergey

      * working on software for inventory database
      * inserted data from Fernando's spreadsheets
      * generating stickers
      * intended for all electronics boards, crates, etc.
      * excludes cables
      * gui for database being developed

Fernando

      * crates will be collected
      * start counter parts here, assembly will start next week
      * pair spectrometer read-out
           + power dissipation and cooling options being considered,
             probably will go with forced convection
           + Co-op student (Ji?) working with Jim Fochtman, designing tray
             to hold SiPMs
      * all TDCs here, Ed upgrading firmware
      * F125s, 98% received, 77% tested
      * cables
           + 4 students working on these, great progress
      * pair spectrometer read-out in progress
           + UConn will install connectors
           + pre-amps installed here, connectors there, back here for SiPM
             installation

Yi

      * solenoid
           + PXS system sends alarm and redundant signals to PLC
           + installed new card to read-out all channels for field mapping
           + inclinometer broke, Scot talking to vendor on options
           + all wiki pages for mapping done
      * goniometer
           + found place to work in FEL
           + FEL looking for support from Hall D
      * microchannel plate development
           + invitation from Argonne see their operation and to talk about
             Hall D
           + may get a seminar here
      * Dry run for China-US conference talk, 4 pm tomorrow
           + USTC, Institute of HEP possible collaborators, strength in
             analysis

    Retrieved from
"https://halldweb1.jlab.org/wiki/index.php/Private:Hall_D_Staff_Meeting,_July_1,_2013"

-- 
Mark M. Ito, Jefferson Lab, marki at jlab.org, (757)269-5295



More information about the Halld-jlab mailing list